OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the points (−3, 2) and (2, 1).

OpenStudy (johnweldon1993):

Point slope form.... first you need to solve for slope given these 2 points.... (-3, 2) (2,1) (x1,y1)(x1,y2) \[m = \frac{ y2 - y1 }{ x2 - x1 }\] That will give you slope Then...pick either one of the points you were given...remember (x,y) and plug THAT point along with the slope you just found into \[y - y1 = m(x - x1)\] take your 'y' value in your point and put it in for 'y1' take your 'x' value in your point and put it in for 'x1' take your slope you just found and put it in for 'm' Then simplify...that will be your point slope formula